Hydrogen iodide undergoes decomposition according
Answer #1

1st calculate initial pressure of HI using
P*V = n*R*T
P*8.00 = 0.316*0.0821*500
P = 1.621 atm

2HI (g) <---------------> H2 (g) + I2 (g)
1.621 0 0 (initial)
1.621-2x x x (at equilibrium)

Kp = x*x/ (1.621-2x)^2
0.060 = x^2/ (1.621-2x)^2
take square root on both sides,
0.245 = x / (1.621-x)
0.397 - 0.245*x = x
x = 0.32 atm

p(I2) = x = 0.32 atm

Answer: e